OpenStudy (anonymous):

Determine which equation below has a solution of a = 2. A. 2a + 3a – 6 = –21 B. 8 = 3 + 4a – 3 – 5a C. –2 = 2 + 5a – 2 – 6a D. –4a + 3a – 6 = –4

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

2 + 5 x 2 - 2 - 6 x 2 = -2.

OpenStudy (anonymous):

better way –2 = 2 + 5a – 2 – 6a -2 = 2 - 2 +5a - 6a -2 = - 1a 2 = a